OpenStudy (youngnephew34):

ABCD is a rectangle. find the length of each diagonal AC= 3y/5 BD= 3y-4

OpenStudy (alexandervonhumboldt2):

ac=bd

OpenStudy (alexandervonhumboldt2):

diagonals of a rectangle are congruent

OpenStudy (alexandervonhumboldt2):

3y/5=3y-4

OpenStudy (alexandervonhumboldt2):

y=5/3

OpenStudy (alexandervonhumboldt2):

now substitute 5/3 into equations

OpenStudy (alexandervonhumboldt2):

ac=5/3*3/5=1=bd bd=1 ac=1

OpenStudy (youngnephew34):

but how?